### v1

Abstract
Short Text Classification (STC) is crucial for processing and comprehending the brief but substantial content prevalent on contemporary digital platforms. The STC encounters difficulties in grasping semantic and syntactic intricacies, an issue that is apparent in traditional pre-trained language models. Although Graph Convolutional Networks enhance performance by integrating external knowledge bases, these methods are limited by the quality and extent of the knowledge applied. Recently, the emergence of Large Language Models (LLMs) and Chain-of-Thought (CoT) has significantly improved the performance of complex reasoning tasks. However, some studies have highlighted the limitations of their application in fundamental NLP tasks. Consequently, this study sought to employ CoT to investigate the capabilities of LLMs in STC tasks. This study introduces Quartet Logic: A Four-Step Reasoning (QLFR) framework. This framework primarily incorporates Syntactic and Semantic Enrichment CoT, effectively decomposing the STC task into four distinct steps: (i) essential concept identification, (ii) common-sense knowledge retrieval, (iii) text rewriting, and (iv) classification. This elicits the inherent knowledge and abilities of LLMs to address the challenges in STC. Surprisingly, we found that QLFR can also improve the performance of smaller models. Therefore, we developed a CoT-Driven Multi-task learning (QLFR-CML) method to facilitate the knowledge transfer from LLMs to smaller models. Extensive experimentation across six short-text benchmarks validated the efficacy of the proposed methods. Notably, QLFR achieved state-of-the-art performance on all datasets, with significant improvements, particularly on the Ohsumed and TagMyNews datasets.
keywords:
[…]
1 Introduction
Short texts are crucial to the contemporary flow of information, particularly given the rapid growth of the Internet[1]. They are essential on major social media platforms, including Twitter, TikTok, Instagram, and Weibo, where they facilitate social interaction and are integral to daily activities. As a vital task for intelligent empowerment and the application of short texts, short-text classification (STC) is essential for applications such as news categorization[2], question answering (QA)[3], and sentiment analysis[4]. Traditional pre-trained language models (PLMs) struggle with the semantic and syntactic intricacies of STC[1, 5]. As shown in Fig 1, categorizing a news title such as ‘Del Potro says make French Open’ into the ‘sport’ category can be challenging, especially without recognizing ‘Del Potro’ as a professional tennis player. The absence of a clear subject and predicate impedes the model’s understanding and learning capacity. To address the challenges posed in STC tasks, Graph Convolutional Networks (GCNs) have shown a certain degree of progress by using an additional knowledge base and redefining STC as a node classification issue, which mitigates the problem of limited training data. However, the effectiveness of GCNs is limited by the quality and extent of their knowledge.
Figure 1: This diagram compares two approaches to the STC task. Due to misinterpretation, the traditional approach erroneously classifies the input ‘Del Potro says make French Open’ as ‘world’. Conversely, our method employs a sequential analytical process that correctly identifies ‘Del Potro’ as a tennis player, recognizes the ‘French Open’ as a tennis tournament, and detects the absence of a grammatical object in the sentence, resulting in accurate categorization under ‘sport’.
The NLP landscape has been revolutionized by Large Language Models (LLMs). These models have performed state-of-the-art downstream tasks such as complex reasoning and QA[6]. Research indicates that models with larger parameters effectively serve as implicit knowledge bases, offering superior integration and application of knowledge compared to traditional external knowledge bases and exhibiting emergent abilities such as in-context learning and instruction following[7, 8]. However, some studies observed limitations in applying LLMs to traditional NLP tasks[9, 10, 11]. For instance, in the named entity recognition task, the performance still lags significantly behind the current best NER model[9]. Additionally, effectively handling datasets for STC tasks, such as Ohsumed and TagMyNews, continues to be challenging. Therefore, our study aimed to leverage the inherent knowledge and emergent capabilities of large models to address this challenge.
A notable approach to enhancing the utilization of knowledge and emergent abilities in LLMs is Chain-of-Thought (CoT) prompting[12]. This technique effectively prompts LLMs to exhibit reasoning processes by necessitating the articulation of intermediate problem-solving steps. Consequently, our study focuses on fostering CoT capabilities within LLMs to facilitate the generation of rationales that directly navigate the intricacies inherent in STC tasks. Furthermore, we seek to foster similar capabilities in smaller language models by transferring knowledge from LLMs, particularly in contexts with limited computational resources or when costly API calls are required to access CoT-capable LLMs.
Specifically, we propose Quartet Logic: A Four-Step Reasoning (QLFR) framework. This framework mainly includes Syntactic and Semantic Enrichment CoT (SSE-CoT), which decomposes the STC task into four more manageable subtasks: (i) key-concept identification, which identifies critical words in the input text; (ii) common-sense knowledge retrieval, which retrieves common-sense knowledge pertinent to these keywords, bridging semantic gaps in short texts; (iii) text rewriting, which rephrases the texts using this knowledge to enhance syntax and readability; and (iv) short-text classification, which employs these refined texts for classification. This strategic design effectively addresses the STC challenges through guided CoTs.
Additionally, we expanded our framework to propose the method CoT-Driven Multi-task learning (QLFR-CML) method, which boosts smaller models by transferring task-aware knowledge from LLMs. This approach aligns closely with practical application requirements for real-world scenarios. In this approach, we extract task-specific rationales from both SSE-CoT, which acquires common-sense knowledge, and Domain Augmentation CoT (DA-CoT), which garners domain knowledge. Subsequently, we employ multi-task learning to fine-tune the smaller models by integrating the ground truth with the rationales from these two distinct CoTs. Moreover, our Explicit Category Context Augmentation (ECCA) strategy refines the model’s output by aligning it more closely with the ground truth.
To sum up, our contributions are as follows:
  1. 1.
    To the best of our knowledge, this research is the first to employ Large Language Models alongside chain-of-thought reasoning to investigate and tackle challenges in short-text classification STC tasks.
  2. 2.
    We introduce Quartet Logic: A Four-Step Reasoning (QLFR) framework incorporating a novel Syntactic and Semantic Enrichment Chain-of-Thought (SSE-CoT) approach. This approach enables LLMs to effectively decompose STC tasks and address semantic sparsity and syntactic ambiguity by breaking them into more manageable subtasks.
  3. 3.
    We propose a CoT-Driven Multi-task learning (QLFR-CML) method designed to enhance the performance of smaller models for STC tasks. This framework leverages task-aware knowledge transfer from LLMs to address the inherent challenges in STC tasks, which can alleviate the challenges of STC tasks in real-world scenarios.
  4. 4.
    Comprehensive experiments were conducted using six challenging short-text benchmark datasets. The experiments confirm that our QLFR framework can significantly utilize LLMs for this challenging task and is superior to several novel baselines. QLFR-CML also shows its capacity to enhance smaller models, even with limited computational resources.
2 Related Work
2.1 Text Classification
Text classification (TC) is a pivotal task in Natural Language Processing NLP that assigns predefined labels to text entities[13]. Historically, TC has been performed through a two-stage process: feature extraction using techniques such as term frequency-inverse document frequency[14], followed by applying classifiers such as support vector machines[15]. Deep learning has revolutionized this approach, allowing for an end-to-end methodology. Contemporary models, including convolutional neural networks[16] and long short-term memory networks[17], can learn independently from raw data, bypass manual feature engineering, and enhance classification adaptability. The introduction of BERT[18] marked a new phase in deep learning and has become a prevalent choice for TC research.
Graph neural networks (GNNs) have emerged as a significant development in TC. Some GNNs, such as TLGNN[19], TextING[20], and HyperGAT[21], treat each document as a network of interconnected word nodes, effectively transforming the text classification challenge into a graph-based task. Although this method introduces a novel viewpoint, its effectiveness can diminish in situations with limited labeled data. On the other hand, models like TextGCN[2] and TensorGCN[22] take a broader approach, framing the classification task within corpus-level graphs where both individual words and entire texts serve as nodes. These models use node classification techniques to identify and classify unlabeled textual elements. However, they may falter when handling concise textual data or data with limited contextual richness.
Recently, the introduction of ChatGPT has revolutionized the field. Numerous studies have used large Language Models for TC tasks. Some studies have leveraged ChatGPT for automated genre recognition to streamline the text classification process through its zero-shot classification abilities[23] and assessed the capacity of ChatGPT for text classification within affective computing by employing it for personality prediction, sentiment analysis, and suicidal ideation detection tasks[24].
2.2 Short Text Classification
Short Text Classification (STC) has consistently emerged as a significant obstacle in the field of NLP, presenting unique and complex challenges in sharp contrast to those encountered in traditional TC tasks[25, 13, 26]. First, the brevity of a text inherently limits its semantic and syntactic integrity[1, 5]. Scholars have sought methods to augment the expressiveness of short texts by integrating additional information[27, 28, 29]. Commonly employed instances include concepts found in external knowledge bases such as [30] and latent topics uncovered within the corpus[31, 32]. Second, STC frequently encounters the dilemma of sparsely labeled data in practical applications[4, 33], thereby increasing the task complexity. A prevailing strategy to mitigate this entails employing graph-based methods, which not only supplement additional information but also offset the paucity of label data, as evidenced by the approach adopted by [34], constructing a corpus-level graph that models latent topics, entities, and documents jointly, where the entities are words linked to knowledge graphs. SHINE[35], which introduces a hierarchically organized heterogeneous corpus-level graph consisting of word-level and document-level graphs, can fully exploit interactions between nodes of the same types and capture similarities between short texts. ST-Text-GCN[36] uses a self-training method for keyword extraction to fully utilize limited labeled texts and a large number of unlabeled texts.
Despite the advancements in graph-based methodologies, they exhibit clear limitations in specific practical scenarios. One of the primary drawbacks is the necessity to retrain the entire model to incorporate new test samples, which can be both computationally intensive and time-consuming. Consequently, there has been a growing interest in inductive reasoning approaches, which obviate the need for retraining by integrating new samples directly with existing training and unlabeled data. Innovations, such as HGAT-inductive[37], propose a novel framework for inductively connecting each new sample to an existing corpus, thereby facilitating dynamic learning. SimpleSTC[38] adopts a word-only approach to navigate the inductive STC problem accurately. Nevertheless, although these inductive methods offer increased flexibility and efficiency, they may necessitate a trade-off in terms of precision because they rely on the assumption that new samples share similar characteristics with the existing corpus.
2.3 Large Language Models
Large Language Models (LLMs), including the ChatGPT[39], LLama2[40], and ChatGLM series[41], have attracted significant attention due to their innovative dialogue system breakthroughs and their potential for various downstream applications. These LLMs have excelled in embedding extensive common knowledge and demonstrating diverse capabilities[7, 8]. The CoT, which enhances LLMs’ reasoning capacities of LLMs by integrating a systematic, stepwise reasoning process within demonstrations, has been applied to various complex reasoning tasks[12, 42, 43].
Presently, research focusing on LLM over the CoT framework principally navigates through two major trajectories. Initially, emphasis was placed on devising sophisticated, human-like cognitive design strategies with methods such as the Tree of Thought[44], which seamlessly linearizes thought processes to facilitate the retrospective exploration and probing of various inferential branches emanating from a central idea. In turn, the Graph of Thought[45] evolves the tree structure into a directed acyclic graph, embedding self-circulation. In contrast, the array of thoughts[46] maintains a single, perpetually evolving context chain, thereby obviating the need for redundant inquiries within the cognitive tree. Subsequently, investigations were conducted to determine the optimal strategies for enhancing the performance of the LLM in assorted tasks. For example, [47] progressively extracted aspects and opinions to navigate the complexities of implicit sentiment analysis, [10] enhanced the relation extraction with heightened precision.
In recent years, the burgeoning scale of deep learning models has been remarkable, with architectures such as transformers and MOE-facilitating models that exceed a trillion parameters, particularly within the era of Large Language Models (LLMs). While beneficial for performance, this escalation in model size has significantly increased deployment costs and impacted inference efficiency, elevating knowledge distillation to a critical role in model compression. Knowledge distillation[48], which is a technique designed to transfer insights from a voluminous teacher model to a more compact student model, has traditionally focused on imparting general knowledge to LLMs and their smaller counterparts, a process exemplified by approaches such as MINILLM[49] and GKD[50]. Moving beyond conventional methods, some strategies leverage the emergent capabilities of LLMs, such as GPT-3[39], which exhibits unique abilities that can be harnessed through in-context learning, chains of thought, and explicit instruction following. Techniques such as the CoT method incorporate intermediate reasoning into prompts, enhancing models under small-sample learning paradigms, such as Meta-Context Tuning and Multitask-Context Tuning, as seen with MT-COT[51]. Furthermore, approaches such as the IF method, as demonstrated by Lion[52], exploit LLMs’ adaptability of LLMs to generate challenging instructions that fortify the capabilities of student models. These innovative methods provide new insights into knowledge distillation and expand the horizons for model compression, marking a significant stride in the efficient utilization of LLMs.
3 Method
In section 3.1, we define the STC task and then reiterate its primary challenges. Section 3.2 introduces Quartet Logic: A Four-step Reasoning Framework, a specialized CoT method designed for LLMs to address the STC task. Finally, we detail the CoT-Driven Multi-task learning method developed to enable smaller models to tackle the STC task effectively in section 3.3.
3.1 Task Description
Given a short text $X_{i}$ belonging to $D$ and a dataset consisting of N short texts, the STC task aims to classify $X_{i}$ as relevant label $l$ from a predefined set of labels $L$.
Semantic and syntactic challenges are inherent in STC tasks owing to their concise and informal nature. State-of-the-art methods based on GCN are constrained by their dependence on an external knowledge base. To address these limitations, we propose distinct frameworks for LLMs and smaller models that target the issues of limited context and reliance on external knowledge.
3.2 Quartet Logic: A Four-Step Reasoning Framework
To improve the performance of LLMs in STC tasks, this study introduces Quartet Logic: A Four-Step Reasoning (QLFR) framework, a CoT-prompt-based approach for multi-step reasoning. Rather than using conventional single-step prompting, which usually yields immediate label predictions without intermediary reasoning stages, we aimed to elicit inherent knowledge and abilities before determining the predicted label.
Our CoT method, called Syntactic and Semantic Enrichment CoT (SSE-CoT), provides a four-step reasoning approach, as shown in Figure 2. The first step involves identifying the key concepts within the input text $X_{i}$. Next, the method retrieves the relevant common sense knowledge $S$ embedded within the LLM. Subsequently, the text was revised to $R$ to enhance the clarity and context, enriching its interpretative depth. Finally, the revised text is used as a query to determine the predicted label, $\hat{y}$.
Figure 2: This diagram presents the Syntactic and Semantic Enrichment CoT (SSE-CoT), as applied to the short text ‘Del Potro says make French Open’. It begins by identifying key concepts, ‘Del Potro’ and the ‘French Open’, then combines them to contextualize ‘Del Potro’ as a tennis player and the ‘French Open’ as a major tournament. The third step refines this information for accuracy and integration. Finally, the process classifies the outcome under ‘sport’. The framework offers a novel solution that effectively addresses STC task challenges.
We conduct our four-step prompts as follows.
Step 1. Key Concept Identification
In the initial step, a specified template is used to query the LLM and pose questions, utilizing its proficiency in identifying relevant concepts. $C^{1}_{1}$[Given the short text $X_{i}$], identify key concepts.
Here, $C^{1}_{1}$ comprises the context of the first step, and the following content constitutes the reasoning instruction $I^{1}_{1}$ for the first step. This stage was designed to focus the model on essential content in preparation for the next steps. The process can be formally expressed as:
$K^{1}=f_{\mathrm{identify}}(C^{1}_{1},I^{1}_{1})$ (1)
where $f_{\mathrm{identify}}$ denotes a function that captures the capability of the models to extract concepts.
Step 2. Common-sense Knowledge Retrieval
With the fundamental concepts $K^{1}$ identified in Step 1, this step involves retrieving the associated common-sense knowledge from the inherent knowledge base of the model using the following template: $C^{1}_{2}$[$C^{1}_{1}$,$K^{1}$], retrieve related common knowledge.
In this phase, concatenate $C^{1}_{1}$ and $K^{1}$ to form the context, and use the following content $I^{1}_{2}$ as the retrieval directive:
$S=f_{\mathrm{retrieve}}(C^{1}_{2},I^{1}_{2})$ (2)
Here, $f_{\mathrm{retrieve}}$ is a function enabling the model to recall pertinent information from an internal knowledge repository. Knowledge retrieval mitigates the semantic gap in STC tasks because of its brevity and facilitates the provision of contextually rich and comprehensive responses.
Step 3. Text Rewriting
Following the retrieval of pertinent common-sense knowledge, this step entails assimilating $S$ into a cohesive and polished short text. The context is formed by concatenating $C^{1}_{2}$ and $S$, where $I^{1}_{3}$ serves as the modification directive. $C^{1}_{3}$[$C^{1}_{2}$,$S$]. Refine and enhance the language to guarantee precision, fluidity, and legibility, whilst preserving the accuracy and wholeness of the integrated information.
Integration, represented by function $g$, is essential for converting raw data into an easily understandable format. It overcomes the syntactic constraints of short texts, producing a structured output $R$ that simplifies comprehension and subsequent categorization by LLM.
$R=g(C^{1}_{3},I^{1}_{3})$ (3)
Step 4. Short Text Classification
Given the short text $R$. classify it into one of the categories. The categories are ‘health’, ‘sport’, ‘entertainment’, ‘business’, ‘sci_tech’, ‘U.S.’ and ‘world’.
After integrating common-sense knowledge and refining the short text $X_{i}$ as $R$, we prompted the LLM with instruction $I^{1}_{4}$ to generate the final predicted label.
$\hat{y_{i}}=\mathrm{argmax}p(y|X^{\prime},I_{4})$ (4)
The label with the highest output probability is designated as the predicted label $\hat{y_{i}}$.
3.3 CoT-Driven Multi-task Learning Method
We propose a novel approach, the CoT-Driven Multi-Task learning (QLFR-CML) framework, specifically designed for smaller models to address STC tasks. The architecture of this framework is illustrated in Figure 3. Our framework consists of two stages, as detailed below.
Figure 3: Overview of the QLFR-CML method. In the first stage, the framework employs SSE-CoT and DA-CoT to prompt LLM with training data for rationale generation. In the second stage, the generated rationales guide the training of a smaller, specialized model. This stage involves multi-task fine-tuning that incorporates a supervised signal, which includes a label and two distinct rationales derived from SSE-CoT and DA-CoT reasoning.
In the initial stage, we employed two specialized CoT prompt processes to generate a rationale from an LLM. The Syntactic and Semantic Enrichment CoT (SSE-CoT) enhances the clarity and coherence of short texts by addressing their syntactic and semantic shortcomings. Simultaneously, Domain Augmentation CoT (DA-CoT) broadens the textual context by integrating domain-specific knowledge.
In the second stage, knowledge transfer occurs from the LLM to a smaller model through a multi-task learning strategy. The smaller model was trained to predict labels and accurately generate common-sense and domain-specific rationales. This integrated training approach improves the reasoning capabilities of the model and strengthens its ability to classify short texts with precision and depth.
We introduce rationale generation in section3.3.1, followed by an overview of Explicit Category Context Augmentation in section 3.3.2, providing a direct and efficient prompt. Finally, we present our multi-task learning strategy in section3.3.3.
3.3.1 Rationale generation
We introduce two distinct CoTs to generate rationales to enhance knowledge transfer in smaller models. First, SSE-CoT, introduced in Section LABEL:s3.2, addresses common-sense knowledge, encompassing broad, universally acknowledged information, such as fundamental life facts and recognized concepts. Second, Domain Augmentation CoT (DA-CoT) focuses on domain-specific knowledge, which includes detailed field-specific information, specialized jargon, and unique concepts. The DA-CoT uses a two-step reasoning process, as shown in Figure 4. We take the output $R$ from the third step of the SSE-CoT and the output $O$ from the second step of the DA-CoT as rationales.
Figure 4: The figure depicts the two-phase procedure of the DA-CoT method employed in the snippets domain. Initially, the method discerns essential text elements, including primary entities, actions, and events. Subsequently, it synthesizes the interconnections and collective importance of these elements, enhancing comprehension of their pertinence and consequences in the context of the text.
We conduct the two-step DA-CoT prompts as follows.
Step 1. Key Concept Identification
In the first step, we used a designated template to query the LLM, capitalizing on its capacity to pinpoint pertinent concepts. Unlike the first step of the SSE-CoT, the DA-CoT incorporates domain-specific cue words. For example, in the news domain illustrated in Figure. 4, it is essential to include critical entities, actions, and events. Details of the other domains are provided in Appendix A. $C^{2}_{1}$ comprises the context of the first step, and the following content constitutes the reasoning instruction $I^{2}_{1}$ for the first step.
$C^{2}_{1}$[Given the short text $X_{i}$], identify the key components, consider the main entities, actions, and events described.
$K^{2}=f_{\mathrm{identify}}(C^{2}_{1},I^{2}_{1})$ (5)
Here, $K^{2}$are the key snippet concepts extracted from $X_{i}$ using $f_{\mathrm{identify}}$ representing the model’s capability to identify and understand key snippet concepts and terminologies from the text.
Step 2. Domain Knowledge Retrieval
After establishing foundational concepts in the initial phase, this step prompts the LLM to apply domain-specific terminology and deeper analytical perspectives to its outputs. $C^{2}_{2}$[$C^{2}_{1}$,$K^{2}$]. Provide a summary of the identified components, including their interrelations and the overall significance within the context of the text.
$O=f_{\mathrm{enrich}}(C^{2}_{2},I^{2}_{2})$ (6)
where $O$ represents the enriched knowledge retrieved or generated by the function $f_{\mathrm{enrich}}$, integrating the identified concepts with in-depth, domain-specific information.
3.3.2 Explicit Category Context Augmentation
In contrast to encoder-only models such as BERT, the smaller model utilized in our study adopts an encoder-decoder architecture. This generative approach requires the prompt enhancement of the comprehension and response of the model to the task. To circumvent the unpredictability inherent in manually crafted prompts, we introduce the Explicit Category Context Augmentation (ECCA) method, which eliminates the requirement for manually crafted task-specific prompts and enriches text representation, leading to a more accurate model classification.
In the ECCA method, the original input text $X_{i}$ is augmented with category labels $L$ using an injection function to form an enhanced input $X_{i}^{\prime}$, which the model subsequently uses for classification. This process was designed to infuse label-specific semantic cues into a model classification task.
$X_{i}^{\prime}=\text{inject}(X_{i},L)$ (7)
The injection can be executed as a simple concatenation, where # represents the concatenation operation and $l_{j},j\in(1,2,\cdots,m)$ denotes each label.
$X_{i}^{\prime}=l_{1}\oplus l_{2}\oplus\ldots\oplus l_{m}\oplus X_{i}$ (8)
3.3.3 Multi-task learning
Several methods exist for integrating rationale into the training processes of downstream models. The direct method uses rationale as an additional input. However, this approach requires that an LLM generate a rationale before the smaller model makes a prediction. Therefore, we adopt a multi-task learning framework to enhance the link between the input $X_{i}$ and desired output $Y_{i}$.
The primary task is to predict the correct category label $\hat{y_{i}}$ from augmented input $X_{i}^{\prime}$. The secondary task involves processing the original text $X_{i}$ to output a rationale $\hat{r_{i}}$, with $r_{i}\in R$ serving as ground truth. Similarly, the tertiary task processes the original text $X_{i}$ to generate the domain-specific rationale $\hat{o_{i}}$, where $o_{i}\in O$ is the ground truth. This multi-task setup aims to predict the category label directly and generate rationales that provide interpretability and context to the model’s decisions, leveraging the rationales from the teacher model as supplementary guidance. Thus, the loss function encompasses the following terms for each task:
$L=L_{label}+\lambda_{1}L_{\text{SSE}}+\lambda_{2}L_{\text{DA}}$ (9)
The weights $\lambda_{1}$ and $\lambda_{2}$ balance the influences of the secondary and tertiary tasks, ensuring that any single task does not dominate the model’s training. The calculation method for each loss function is as follows:
$L_{label}=\frac{1}{N}\sum_{i=1}^{N}\ell(f_{s}(X_{i}^{\prime}),\hat{y_{i}})$ (10)
$L_{SSE}=\frac{1}{N}\sum_{i=1}^{N}\ell(f_{s}(X_{i}),r_{i})$ (11)
$L_{DA}=\frac{1}{N}\sum_{i=1}^{N}\ell(f_{s}(X_{i}),o_{i})$ (12)
Here, $f_{s}$ represents the smaller model, and $\ell$ denotes the cross-entropy loss between the predicted and target tokens. The estimated rationales are unnecessary during testing, thereby obviating the need for LLM at that stage.
4 Experiment
4.1 Datasets
To ensure a thorough and unbiased evaluation, we conducted extensive experiments on six widely recognized benchmark short-text datasets: MR, Snippets, Ohsumed, StackOverflow, TagMyNews, and AGNews. Following [35], we randomly sampled 40 labeled short texts from each class, where half formed the training set, and the other half formed the validation set. Table 1 provides detailed information regarding the datasets. We describe the datasets further below.
#texts avg.length #classes #train(radio)
MR 10662 12.1 2 40(0.38%)
Snippets 12340 17.4 8 160(1.30%)
Ohsumed 7400 8.65 23 460(6.22%)
StackOverflow 20000 5.7 20 400(2.00%)
TagMyNews 32605 6.1 7 140(0.43%)
AGNews 20000 27.8 4 80(0.4%)
Table 1: Summary of short text datasets used.
  1. 1.
    MR [^1] https://www.cs.cornell.edu/people/pabo/movie-review-data/: introduced by [53], consists of one-sentence movie reviews annotated as either positive or negative for sentiment analysis.
  2. 2.
    Snippets [^2] https://github.com/jacoxu/STC2/tree/master: introduced by [1], are snippets derived from Google search engine results.
  3. 3.
    Ohsumed [^3] https://github.com/yao8839836/text_gcn: introduced by [54], this dataset focused on classifying cardiovascular diseases. In this study, we utilized the subset defined by [34], which concentrates on the classification of short texts using only the titles of single-label documents.
  4. 4.
    StackOverflow [^4] https://github.com/jacoxu/StackOverflow/blob/master/: introduced by [2], contains 20 distinct types of question titles, derived from the StackOverflow data within the IT Q&A community.
  5. 5.
    TagMyNews [^5] https://github.com/cskarthik93/TagMyNews-Classification: introduced by [34], comprises English news headlines sourced from the Really Simple Syndication (RSS) feeds.
  6. 6.
    AGNews [^6] https://github.com/mhjabreel/CharCNN: introduced by [55], was the source of our dataset. In this study, we randomly selected 20,000 articles from the collection.
4.2 Experimental Setting
We selected LLaMA2-13B[40] as the foundational model for our QLFR method and chose the LLaMA2-13B and Flan-T5-Large[56] models to represent the LLM and smaller model in the QLFR-CML method, respectively. In the QLFR method, the LLaMA2-13B employs efficient parameter fine-tuning with Low-Rank Adaptation (LoRA)[57], which is conducted with a batch size of 10 across five epochs. LoRA maintains the weights of pretrained LMs while introducing trainable rank decomposition matrices into each transformer layer, making it feasible to fine-tune larger LMs with fewer computational resources [^7] In our experiment, trainable parameters only account for 0.24% of the entire LLaMA2-13B parameters. Conversely, in the QLFR-CML, Flan-T5-Large has fully fine-tuned parameters with batch sizes of 5 and 10 epochs. All experiments were conducted using three A100s and five V100s.
4.3 Evaluation
To assess the efficacy of the proposed approach, we selected two principal evaluation metrics: accuracy (denoted as ACC) and macro-averaged F1 score (denoted as F1). Accuracy provides a straightforward measure of the overall accuracy of the model predictions. Conversely, the macro-averaged F1 score is crucial for imbalanced datasets because it uniformly weighs precision and recalls across all classes, ensuring fair evaluation even with limited samples in some categories.
4.4 Compared Methods
The baselines can be divided into three primary categories: group (A), pre-trained language models; group (B), GCN-based Models; and group (C), Large Language Models. Each group is discussed below.
Group (A). Pre-trained Language Models
Pre-trained Language Models (PLMs) have garnered considerable attention in the field of NLP, frequently serving as tools for text classification, among other basic tasks. In this study, we chose three established methods for comparative analysis.
  1. 1.
    BERT [^8] https://tfhub.dev/tensorflow/bert_en_uncased_L-12_H-768_A-12/4[18], pre-trained on extensive corpora, is further fine-tuned using a linear classifier for short-text classification. Each document can be represented by either the average of its word embeddings (denoted by -avg) or the embedding of a CLS token (denoted by -CLS).
  2. 2.
    RoBERTa [^9] https://github.com/facebookresearch/fairseq[58] is a modified and optimized version of BERT that is pre-trained on large amounts of text using adjusted training approaches.
Group (B). GCN-based models
Recently, methods based on Graph Convolutional Neural Networks (GCNs) have achieved outstanding results in STC tasks. Five classical models are used in this study.
  1. 1.
    HGAT-inductive [^10] https://github.com/BUPT-GAMMA/HGAT[37] advocates connecting each new sample to the existing training and unlabeled data in the corpus.
  2. 2.
    SimpleSTC [^11] https://github.com/tata1661/SimpleSTC-EMNLP22[38] constructs a basic word graph of common words using an additional corpus to address STC tasks via inductive learning.
  3. 3.
    ST-Text-GCN [^12] https://github.com/wanggangkun/ST-Text-GCN[36] utilizes a self-training method to extract keywords, ensuring the effective use of limited labeled text and many unlabeled texts.
  4. 4.
    HGAT [^13] https://github.com/BUPT-GAMMA/HGAT[34] deploys a dual-level attention-based GNN to function on a corpus-level graph encompassing entities, topics, and documents.
  5. 5.
    SHINE [^14] https://github.com/tata1661/SHINE-EMNLP21[35] introduces a hierarchically heterogeneous corpus-level graph that optimizes node interactions and captures textual similarities.
Group (C). Large Language Models
Four LLMs were selected, with LLaMA2-7B and ChatGLM used for comparative experiments and the GPT-3 and FLAN-T5 series used in subsequent analytical experiments.
  1. 1.
    LLaMA2-7B [^15] https://github.com/facebookresearch/llama-recipes/[40], introduced by Mata, encompasses seven billion parameters. It exhibits superior efficacy on multiple benchmark datasets and is suitable for research and commercial applications.
  2. 2.
    ChatGLM [^16] https://github.com/THUDM/ChatGLM-6B, introduced by Tsinghua University, is a robust language-generation model that provides advanced deep-learning technologies with training using extensive corpora.
  3. 3.
    GPT-3[59], note that GPT-3 does not release the model parameters, and we use them via the API. Consequently, the supervised fine-tuning of GPT-3 is not feasible; it is utilized solely for the experiments described in Section 4.7.
  4. 4.
    FLAN-T5 series [^17] https://huggingface.co/google/flan-t5-xxl[56], introduced by Google. This method fine-tunes language models for tasks of an unprecedented scale owing to the remarkable generalization capacities of these models. Consequently, a singular model can effectively execute more than 1,000 tasks. In Section 4.9, we perform comparative experiments using models of various sizes.
4.5 Benchmark Comparison
MR Snipptes Ohsumed TagMyNews StackOverflow AGNews
ACC F1 ACC F1 ACC F1 ACC F1 ACC F1 ACC F1
PLMs BERT-avg 51.69 50.65 79.31 78.47 23.91 4.98 55.13 44.26 $72.91^{*}$ $73.69^{*}$ $76.52^{*}$ $76.49^{*}$
BERT-CLS 53.48 46.99 81.53 79.03 21.76 4.81 58.17 41.04 $73.74^{*}$ $74.11^{*}$ $78.35^{*}$ $78.42^{*}$
RoBERTa $53.62^{*}$ $52,27^{*}$ $79.58^{*}$ $79.10^{*}$ $26.95^{*}$ $19.47^{*}$ $55.57^{*}$ $50.45^{*}$ $64.87^{*}$ $64.40^{*}$ $79.33^{*}$ $79.45^{*}$
GCNs HGAT-inductive 61.18 59.77 79.40 77.69 42.08 25.71 58.20 49.55 $72.31^{*}$ $70.42^{*}$ 70.23 68.43
SimpleSTC 62.27 62.14 80.96 80.56 $43.16^{*}$ $23.35^{*}$ 67.17 63.34 $73.63^{*}$ $73.39^{*}$ $72.62^{*}$ $71.89^{*}$
ST-Text-GCN $50.23^{*}$ $34.02^{*}$ $83.83^{*}$ $83.15^{*}$ $33.64^{*}$ $22.62^{*}$ $52.33^{*}$ $47.38^{*}$ $69.68^{*}$ $68.94^{*}$ $86.83^{*}$ $86.06^{*}$
HGAT 62.75 62.36 82.36 74.44 42.68 24.82 61.72 53.81 $75.29^{*}$ $75.14^{*}$ 72.10 71.94
SHINE 64.58 63.89 82.39 81.62 45.57 30.98 62.50 56.21 $76.81^{*}$ $76.44^{*}$ $81.39^{*}$ $81.45^{*}$
LLMs LLaMA2-7B $71.49^{*}$ $71.03^{*}$ $78.47^{*}$ $78.76^{*}$ $48.08^{*}$ $40.21^{*}$ $56.75^{*}$ $56.20^{*}$ $87.36^{*}$ $88.21^{*}$ $79.49^{*}$ $80.67^{*}$
ChatGLM $73.50^{*}$ $73.31^{*}$ $80.62^{*}$ $80.39^{*}$ $51.28^{*}$ $36.68^{*}$ $70.77^{*}$ $67.79^{*}$ $86.13^{*}$ $86.77^{*}$ $82.51^{*}$ $82.50^{*}$
Ours QLFR-CML 76.10 76.10 84.67 83.89 58.44 44.32 79.61 75.43 84.11 84.21 88.30 88.41
$\uparrow$ 2.60 $\uparrow$ 2.88 $\uparrow$ 0.84 $\uparrow$ 0.74 $\uparrow$ 7.16 $\uparrow$ 4.11 $\uparrow$ 8.84 $\uparrow$7.64 $\downarrow$ 3.25 $\downarrow$ 4.00 $\uparrow$ 1.47 $\uparrow$ 2.35
QLFR 81.70 81.72 85.75 85.32 61.10 51.85 83.37 79.84 89.67 89.58 89.14 89.28
$\uparrow$ 8.20 $\uparrow$8.41 $\uparrow$1.92 $\uparrow$2.17 $\uparrow$ 9.82 $\uparrow$ 11.64 $\uparrow$ 12.60 $\uparrow$12.05 $\uparrow$ 2.31 $\uparrow$ 1.37 $\uparrow$ 2.31 $\uparrow$ 3.22
Table 2: Performance evaluation measured on short text datasets. This table presents the comparative performance of baseline models and ours, measured in terms of ACC (%) and F1 (%). The table highlights the highest scores in bold, with underscores indicating the highest scores achieved by prior methods. *** indicates that the result is reproduced by us.
Table 2 shows the performance comparison. As can be seen, our QLFR method surpasses other approaches on all six datasets. Similarly, the QLFR-CML method shows notable performance across five datasets. Specifically, on the TagMyNews dataset, QLFR achieves an increase of 12.60% in ACC and 12.05% in F1 score compared to the previously optimal ChatGLM model. QLFR-CML records an 8.84% improvement in ACC and a 7.64% increase in F1 score. These results support our hypothesis that LLMs can effectively utilize their inherent knowledge and abilities to address traditional NLP tasks, particularly the STC task discussed in this paper, thereby validating the effectiveness of our proposed methods.
A comparative analysis revealed that QLFR outperformed QLFR-CML, especially on the MR dataset, where QLFR’s accuracy of QLFR exceeded that of QLFR-CML by 5.6%. This indicates that models with larger parameters have greater intrinsic knowledge and enhanced capabilities to solve the STC task. Although the CML did not outperform ChatGLM on the StackOverflow dataset, it significantly exceeded the top-performing model in the GCN-based group. Moreover, it achieved a 7.3% higher ACC and a 7.77% greater F1 score than SHINE.
An interesting observation is that for the Snippets and AGNews datasets, the methods based on LLMs did not outperform those utilizing GCNs. In the Snippets dataset, ST-Text-GCN outperformed ChatGLM by 3.21% in ACC and 2.76% in F1 score. Given the dense entity relationship characteristics of news datasets, the inherent topological advantages of GCN methods can enable a more effective capture of relational data, leading to superior performance. Consequently, traditional approaches have retained their relevance in the context of LLMs.
4.6 Ablation Study
MR Snipptes Ohsumed TagMyNews StackOverflow AGNews
ACC F1 ACC F1 ACC F1 ACC F1 ACC F1 ACC F1
QLFR-CML 76.10 76.10 84.67 83.89 58.44 44.32 79.61 75.43 84.11 84.21 88.30 88.41
w/o ECCA 75.48 75.46 83.27 81.80 56.91 48.65 77.07 74.06 83.08 83.44 87.52 87.74
w/o SSE-CoT 74.36 74.38 81.55 80.64 56.43 47.96 75.96 72.19 82.87 83.58 86.38 86.80
w/o DA-CoT 74.02 74.02 81.21 80.17 55.18 47.27 75.45 71.73 81.96 82.74 85.87 86.10
Table 3: Performance evaluation of QLFR-CML and its variant on short text datasets. The best results are in bold. ‘w/o ECCA’ illustrates results without employing the ECCA strategy, while ‘w/o SSE-CoT’ and ‘w/o DA-CoT’ display outcomes when omitting SSE-CoT and DA-CoT rationales, respectively.
QLFR-CML Ablation. Ablation studies were performed on six benchmark datasets to assess the influence of particular strategies or components on the QLFR-CML method. We developed three variations of the QLFR-CML, as follows:
  • QLFR-CML+w/o ECCA: During the fine-tuning stage, the original text $X$ is used as input to predict the label without employing the ECCA strategy.
  • QLFR-CML+w/o SSE-CoT: The task of generate SSE-CoT rationales is removed from the multi-task learning process.
  • QLFR-CML+w/o DA-CoT: Similarly, the task of generate DA-CoT rationales is excluded from the multi-task learning process.
Table 3 displays the outcomes of our ablation study, highlighting the best-performing metrics in bold. These findings demonstrate the integral role of each strategy and component in the effectiveness of our QLFR-CML method. Notably, ECCA contributes most substantially, with SSE-CoT and DA-CoT also providing significant enhancements. The significant improvement provided by ECCA confirms the importance of a prompt for generative models with an encoder-decoder architecture. The improvement observed with SSE-CoT and DA-CoT affirms that our strategy effectively transfers the knowledge and capabilities of LLM to a smaller model.
The impact of SSE-CoT and DA-CoT varied according to the dataset type. For instance, in news-related datasets, such as TagMyNews, SSE-CoT exerts a more substantial influence than DA-CoT, likely because of the frequent occurrence of widely recognized entities in news texts. Conversely, in specialized datasets, such as Ohsumed for medical content and StackOverflow for computing science, the importance of DA-CoT increases, reflecting the prevalence of domain-specific terminology.
MR Snipptes Ohsumed TagMyNews StackOverflow AGNews
ACC F1 ACC F1 ACC F1 ACC F1 ACC F1 ACC F1
QLFR 81.70 81.72 85.75 85.32 61.10 51.85 83.37 79.84 89.67 89.58 89.14 89.28
w/o rewriting 81.38 81.38 85.28 84.16 60.80 49.63 82.78 78.91 89.08 88.44 88.52 89.04
w/o retrieval 80.18 80.20 83.65 83.72 57.43 42.96 81.04 76.48 87.87 87.58 87.30 88.61
w/o both 79.84 79.84 83.16 83.51 56.99 40.65 80.44 76.12 87.35 87.51 87.02 87.94
Table 4: Performance evaluation of QLFR and its variant on short text datasets. The best results are in bold. ‘w/o rewriting’ and ‘w/o retrieval’ refer to results without the rewriting step and retrieval step in SSE-CoT, respectively. ‘ w/o both’ refers to the exclusion of both steps
QLFQ Ablation. Ablation studies were performed on six benchmark datasets to assess the influence of the components of the QLFQ framework. We developed two variations of the QLFQ as follows.
  • QLFR+w/o rewriting: Omission of the rewriting step, using the concatenation of the original input with the retrieved text as input.
  • QLFR+w/o retrieval: Omission of the retrieval step, employing rewritten original inputs as input.
  • QLFR+w/o both: Utilizing the original input without modification.
Table 4 displays the outcomes of our ablation study, highlighting the best-performing metrics in bold. The findings indicate that both steps in our SSE-CoT are beneficial. The rewriting step addresses the challenge of syntactic inexactitude in short texts, while the retrieval step effectively resolves the issue of semantic sparsity. A comparison reveals that retrieval offers slightly more advantage than rewriting, suggesting that semantic deficiencies in short texts are more critical than syntactic imprecision.
4.7 Analysis of In-Context Learning for QLFR
In addition to the supervised fine-tuning (SFT) paradigm, in-context learning has gained popularity. We chose four LLMs, ChatGLM, LLaMA2-7B, LLaMA2-13B, and GPT-3, for experimentation in the zero-shot and one-shot settings. Given the rigorous demands of in-context learning, the QLFR was tested solely on the GPT-3. We manually selected a sample from each category in a one-shot setting and constructed the SSE-CoT as the context. Specific examples of the input formats are provided in Appendix B. The experimental results are presented in Figure 5.
Figure 5: Performance evaluation of LLMs in zero-shot and one-shot settings is conducted using three representative datasets. The upper three groups correspond to zero-shot settings, while the lower three pertain to one-shot settings. In each figure, the best results are highlighted in bold.
Zero-shot setting. In the analysis across the three datasets under a zero-shot setting, the in-context learning abilities of the four models were ranked as follows: GPT-3 outperformed LLaMA2-13B, which, in turn, surpassed both ChatGLM and LLaMA2-7B. This hierarchy is attributable to in-context learning aptitude inherent in larger models, with increased parameters correlating with enhanced performance. Applying our QLFR method to GPT-3 resulted in performance gains across the board. For instance, the ACC on the Ohsumed dataset improved from 51.6% to 52.9%. These findings indicate that our approach is not limited to the SFT but is also applicable to the context learning paradigm.
One-shot setting. In the one-shot setting, the findings regarding the model performance were consistent with those observed in the zero-shot setting. Furthermore, our method achieved improvements across all datasets.
Comparing the results of the zero-shot and one-shot tasks, contrary to our initial expectations, the hypothesis that providing task-relevant examples enhances the comprehension and task response of LLM is not supported. The findings reveal that only the GPT improves performance in one-shot learning compared with zero-shot learning in the two datasets. In contrast, a decline is observed in the remaining datasets. Further analysis of the model outputs suggested that LLM with approximately 10 billion parameters tended to demonstrate a reduced capacity for processing instructions with greater input complexity. In contrast, the GPT consistently exhibits a robust understanding of instructions and sustains its performance despite increased context length.
Comparison of the results of SFT and in-context learning paradigm. GPT-3, with over a hundred billion parameters, exhibits a remarkable in-context learning performance. In the one-shot setting, GPT-3 exceeded QLFR-CML, and GPT-3+QLFR nearly matched the supervised QLFR method. However, for models with billions of parameters, the SFT paradigm remained predominant.
4.8 Analysis of different base models for QLFR-CML
In our QLFR-CML method, the initial stage involves the extraction of rationales with the assistance of an LLM, followed by a fine-tuning phase that requires a smaller model. Therefore, we investigated the outcomes of employing the same LLM with various smaller models and the results of utilizing the same smaller model with different LLMs.
Figure 6: Evaluation of performance across various smaller model sizes reveals that as the parameter count in these models escalates, there is a performance improvement, albeit not in direct proportion.
Different smaller models. As shown in Figure 6, we utilized LLaMA2-13B as the LLM and three versions of Flan-T5-Base, Flan-T5-Large, and Flan-T5-XL as smaller models, with parameter counts of 250M, 780M, and 3 B, respectively. It is apparent that, with an increase in the parameter size of the smaller models, there is a corresponding enhancement in both the ACC and F1 metrics, which suggests a correlation between the capacity of the smaller model and its performance. However, this relationship is not strictly linear. For instance, in the Ohsumed dataset, the increase in ACC from the 250M model to the 780M model does not reflect the expected proportional improvement, indicating diminishing returns as the model size increases.
MR Ohsumed TagMyNews
ACC F1 ACC F1 ACC F1
ChatGLM 70.17 70.54 55.11 47.25 73.75 70.48
LLaMA2-7B 73.89 73.89 55.90 49.03 77.63 74.09
Flan-T5-XXL 72.44 72.65 51.72 44.10 76.71 72.97
LLaMA2-13B 76.10 76.10 58.44 44.32 79.61 75.43
GPT-3 79.46 79.46 59.51 46.85 81.07 77.95
Table 5: Evaluation of performance across various-sized Large Language Models: Larger models demonstrate enhanced knowledge transfer capabilities attributable to their expanded parameter count.
Different LLMs. We selected five models as representatives of LLMs and Flan-T5-Large as a comparatively smaller model. The experimental results presented in Table 5 indicate that GPT-3, as an LLM, transfers knowledge and capabilities most effectively to Flan-T5-Large. On the MR dataset, GPT-3’s ACC and F1 scores exceed those of LLaMA2-13B by 3.36% and 3.36%, respectively. Despite having fewer parameters than Flan-T5-XXL, LLaMA2-7B outperformed it on the Ohsumed and TagMyNews datasets, which may be attributed to differences in the model’s training corpus and strategies.
4.9 Analysis of different prompts for QLFR-CML
To verify the efficacy of the proposed ECCA, we conducted comparative experiments using two prompts with different semantic richness. Details of the prompts are provided in the AppendixC.
MR Ohsumed TagMyNews
ACC F1 ACC F1 ACC F1
w/o prompt 75.48 75.46 56.91 48.65 77.07 74.06
ours 76.10 76.10 58.44 44.32 79.01 75.43
prompt1 75.29 75.30 57.28 48.33 77.62 74.68
prompt2 76.15 76.15 58.26 48.71 78.89 75.36
Table 6: Evaluation of Different Prompts in QLFR-CML: the efficacy of ECCA demonstrated through significant outcomes.
Table 6 lists several interesting trends. Prompts enhanced performance across all measured metrics relative to their absence. Furthermore, our method outperformed the less semantically rich prompt1 and demonstrated greater accuracy than prompt2 in the two datasets. Although our approach does not outperform prompt2 on the MR dataset in terms of ACC, it offers considerable benefits. It obviates the necessity for creating manual, dataset-specific prompts, which yield notably shorter input lengths than prompt2 by simply appending label names, benefiting processing efficiency and model scalability.
4.10 Analysis of training size
We investigated the impact of the training data ratio on the model performance. Experiments were conducted using SHINE, ChatGLM, and our proposed methods on the MR and TagMyNews datasets.
Figure 7: Performance evaluation of different ratios of training data.
The results presented in Figure 7 show that the proposed method significantly outperforms the other two models in low-resource environments. The data revealed a consistent increase in test accuracy correlated with the rising proportion of training data for both the MR and TagMyNews datasets, consistent with supervised learning principles. Additionally, the acceleration or plateau of the ACC improvement suggests diminishing returns in the model performance beyond a certain data threshold. Notably, the ACC of SHINE and ChatGLM approach parity, whereas our methods consistently outperform them, indicating that, with adequate data, the inherent strengths of the model become increasingly influential.
5 Conclusion
In this study, we introduce a novel reasoning framework, Quartet Logic: a Four-Step Reasoning (QLFR) framework, alongside the CoT-Driven Multi-task learning (QLFR-CML) method, to address the challenge of short text classification. The QLFR framework employs large Language Models (LLMs) for resolution, whereas the QLFR-CML method uses smaller models. Our framework leverages extensive, inherent knowledge and diverse abilities that surpass prior GCN-based models. To harness this knowledge within LLMs, we implemented syntactic and semantic enrichment CoTs to facilitate methodological reasoning by LLMs. Moreover, to impart this knowledge to the smaller models, we adopted a multi-task learning approach. We introduced an additional Domain-Adaptive CoT to enrich the knowledge transfer. In the fine-tuning phase for smaller models, we proposed an Explicit Category Context Augmentation mechanism to improve task comprehension. Comprehensive experiments were conducted across six prevalent datasets to determine the effectiveness of the proposed framework. The experimental results demonstrated that the proposed framework significantly outperformed established baselines. However, this complex task remains unresolved because its performance on domain-specific datasets remains suboptimal. Future efforts should focus on integrating LLMs and additional knowledge to refine the proposed approach.
6 Limitations
Our study involved carefully designed experiments to validate the effectiveness of the proposed method. However, limited hardware resources and time constraints affect the ideal number of LLMs employed. Moreover, swift progress in large-language-model technologies poses the challenge of maintaining up-to-date comparative analyses. Nonetheless, we endeavored to ensure the robustness and relevance of our findings within these limitations.
Appendix A DA-CoT: Cross-Domain Applications
Figure 8: DA-CoT of medical domain
Figure 9: DA-CoT of computer science domain
Appendix B In-Context Learning input apllications
Figure 10: zero-shot setting
Figure 11: one-shot setting
Appendix C Different Prompts for QLFR-CML
Categorize this text: ‘wal-mart buys social media firm kosmix’.
Given the short text ‘wal-mart buys social media firm kosmix’, classify it into one of the categories. The categories are health, sport, entertainment, business, sci_tech, U.S. and world.